OpenStudy (anonymous):

The quotient of a number and 2 is the same as the difference of the number doubled and 3. What is the number? 2 -2 -4 1/2

OpenStudy (anonymous):

x/2=2x-3 can you do it from here?

OpenStudy (anonymous):

sorry x/2=3-2x

OpenStudy (anonymous):

but yes it is -2

OpenStudy (anonymous):

k

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

yes